OpenStudy (anonymous):

The numerator of a fraction is 3 times the denominator. If 5 is subtracted from the numerator, and added to the denominator, the resulting fraction is equal to 16/12. Find the original fraction.

OpenStudy (anonymous):

2

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\begin{cases}n=3d\\ \\ \frac{n-5}{d+5}=\frac{16}{12}\end{cases}\]

OpenStudy (anonymous):